OpenStudy (anonymous):

(2x)/(x-5) - x/(x-4) Please simplify

OpenStudy (anonymous):

x(x-3)/(x-4)(x-5) :)

OpenStudy (anonymous):

\[\frac{ 2x }{ x-5}-\frac{ x }{ x-4}\]

OpenStudy (anonymous):

\[\frac{ (2x)(x-4)-x(x-5) }{ (x-5)(x-4) }=\frac{ x ^{2}-3x }{ (x-5)(x-4) }\] \[\frac{ x(x-3) }{ (x-5)(x-4)}\]